OpenStudy (anonymous):

Find the line that passes through(1,2) and the parallel to the line with x-intercept -3 and y-intercept of 10

OpenStudy (zzr0ck3r):

a line with x intercept (-3,0) and y intercept (0,10) has slope \(?\)

OpenStudy (zzr0ck3r):

dont wait for dan to show off, ANSWER ME.....

OpenStudy (zzr0ck3r):

jk

OpenStudy (dan815):



OpenStudy (dan815):

lool

OpenStudy (dan815):

Do it!! find the slope be a man!

OpenStudy (anonymous):

would i use m=y-y1/x-x1

OpenStudy (anonymous):

10-2/-3-1=-2

OpenStudy (zzr0ck3r):

ok yes we would use \(m=\frac{y_0-y_1}{x_0-x_1}\) but our given points are \((-3,0)\) and \((0,10)\)

OpenStudy (zzr0ck3r):

so our slope is ?

OpenStudy (anonymous):

For two lines to be parallel, both lines MUST have the same slope. To find the slope of the given equation, modify it into the "slope-intercept" form of: y = mx + b where m is the slope b is the y-intercept at (0,b) . 5 = -10x - 10y 10y + 5 = -10x 10y = -10x - 5 y = -x - 5 . From the above we know the slope is a -1. Now to find the 'y-intercept' of the new line, use that fact AND the given point of (-2, 1) to plug it back into: y = mx + b 1 = (-1)(-2) + b 1 = 2 + b -1= b . Recapping, we have: m = -1 b = -1 therefore the equation is: y = mx + b y = -x -1 (this is the equation)

OpenStudy (zzr0ck3r):

yikes our line will have slope \(m=\frac{10-0}{0-(-3)}=\frac{10}{3}\)

OpenStudy (zzr0ck3r):

do you see how I got that?

OpenStudy (anonymous):

that's what i got and then i would plug it in y-y1=m(x-x1)